Using only 32-cent and 20 cent stamps, Charlie put $3.36 postage on a package he sent to his sister. He used twice as many 32-cent stamps as 20-cent stamps. Determine how many of each type of stamps he used.

Found 2 solutions by brucewill, richwmiller:
Answer by brucewill(101)

Let x = Number of 20c stamps
Therefore, 2x = Number of 32c stamps
20(x) + 32(2x) = 336
20x + 64x = 336
84x = 336
x = 4 (20c stamps)
x2 = 8 (32c stamps).
Verify:
4 x 20 = 80 cents
8 x 32 = 256 cents
80+256 = $3.36.

Answer by richwmiller(17219)

32x+20y=336
4*(8x+5y) = 336
8x+5y = 84
x=2y
y=4
x=8
check
32*8+20*4=336
256+80=336
check